OpenStudy (anonymous):

Help writing equations using point-slope form?... find the slope and standard form of the line that passes through the points (3,6) and (4,5) help!

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1) then use Point-slope form

OpenStudy (anonymous):

x2 is not x^2.......

OpenStudy (anonymous):

slope between two points is \[m=\frac{ \Delta y }{ \Delta x }\] point slope form: \[y-y _{1} = m (x-x _{1})\]

OpenStudy (anonymous):

One sec..

OpenStudy (anonymous):

Can you explain it better please?

OpenStudy (anonymous):

Like how to plug it in and everything.

OpenStudy (anonymous):

\[m=\frac{ \Delta y }{ \Delta x } = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\] from your points (3,6) and (4,5). Choose one to be point 1 and point 2 (it doesn't matter which one you choose to be which) \[P _{1} = (x _{1},y _{1}), P _{2} = (x _{2},y _{2})\] then plug the point into the equation above.

OpenStudy (anonymous):

I STILL DONT GET IT

OpenStudy (anonymous):

and im getting all of the answers wrong

OpenStudy (anonymous):

let's say \[P _{1}=(3,6), P_2=(4,5)\] plugging this in we get \[m=\frac{ \Delta y }{ \Delta x } = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }= \frac{ 6-5 }{ 3-4 } = \frac{ 1 }{ -1 }=-1\] then plugging that into the equation for y \[y-y _{1} = m (x-x _{1})\] \[y-6 = -1 (x-3)\]

OpenStudy (anonymous):

you must be wrong because your answer isnt in the choices :|

OpenStudy (anonymous):

@xon1300

OpenStudy (anonymous):

the choices all start with m

OpenStudy (anonymous):

slope m = -1, \[y= -x +9\] or \[x+y=9\]